216 small drops of liquid coalesce to form a larger drop. How does the capacity of the drop change?

Grade:12

1 Answers

Eshan
5 years ago
The total volume of the drops before and after coalescence will be same.

Hence216\times \dfrac{4}{3}\pi r^3=\dfrac{4}{3}\pi R^3
\implies R=6r

The capacity of a charged sphere is4\pi\epsilon_0 R\propto R

Therefore the capacity of the drop is 6 times of that of each smaller drop.